# Dominik Bojko Krzysztof Grining Marek Klonowski

### v6

@@ 8.1 Main Theorem for Morris Counter (Theorem 1)
    $\delta$ phase (Theorem 5),
@@
    extrapolation of $\mathcal{P}_{k}^{(c)}\leq 2^{c+3}\mathcal{P}_{k}^{(c+1)}$ property to $N>n_{k}$ (Lemma 3),
@@
    $\varepsilon$ phase (Theorem 6).
During the first phase, we consider a concentration of Morris Counter in the vicinity of its mean value. More precisely, we show that the Morris Counter after $n$ increment requests takes values in relatively small intervals $I_{n}$ with probability $1-\delta$ (note that then $M_{n}$ satisfies condition (2) for $S_{n}=I_{n}$), where $I_{n}$ is defined as in (4) and $\delta$ is some small constant, which arises from the proof. Note that $I_{n}$ can be interpreted as confidence intervals at level $1-\delta$ (see, e.g. [40]). This phase is divided into lemmas 1 and 2. The first one uniformly bounds the formula for probabilities given by the Theorem below, due to Flajolet:
Theorem 4 (Proposition 1 from [24])
@@
In Lemma 1, we sum up the bounds on $p_{n,l}$ to obtain a small upper bound for $\delta_{1}:=\mathbb{P}\left(M_{n}\leqslant\left\lceil\log(n)\right\rceil-5\right)$. The same bounds cannot be utilised efficiently in the proof of Lemma 2. Instead, it couples $M_{n}$ with a process $X_{n}$, which increases during the first $\left\lceil\log(n)\right\rceil+1$ steps, and then follows the same update rule, so $M_{n}\leqslant X_{n}$ almost surely.
Therefore $\delta_{2}:=\mathbb{P}\left(M_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)\leqslant\mathbb{P}\left(X_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)$, which is much easier to bound from definition. Note that such a coupling cannot be used in the proof of Lemma 1. The first phase is summarised by Theorem 5, i.e., establishes $\delta=\delta_{1}+\delta_{2}$.
In the second phase, we show that $(\mathcal{P}_{k}^{(4)})_{k}$ is descending for large enough $k$ and $(\mathcal{P}_{k}^{(5)})_{k}$ is ascending for big enough $k$ (Lemma 6). A change of monotonicity is the first obstacle in the proof of the main theorem. The main idea beneath the proof of this fact is to calculate the differences between consecutive elements of the considered ”special” sequences by representing them as the sums via application of Flajolet’s Theorem 4 and realising that usually at most first ten terms of the sums are crucial (on the other hand, taking less than eight terms is rarely sufficient). This is the second issue, which makes the proof so complicated. Let us note that Theorem 4 presents an explicit formula for $\mathbb{P}\left(M_{n}=l\right)$, which (as we may experience in A) is not convenient to analyse. However, it is simple enough to find the values numerically (also note that recursive Definition 3 provides those probabilities easily as well, However, this approach is inefficient in terms of memory and time for a large number of requests $n$). Therefore, by precise analysis, we can finally check some sums numerically and obtain the thesis of Lemma 6 for $k\geqslant 15$. However, numerically, one can extrapolate it to some smaller $k$ as well (remark that this proof does not work for small values of $k$, since not only ten terms of the aforementioned sums are important). The lemma 6 can be used directly to show that $\mathcal{P}_{k}^{(4)}\leqslant 2^{7}\mathcal{P}_{k}^{(5)}$ for $k\geqslant 7$ (Claim 1). Let us note that, in fact, this result is not true for $k<7$.
@@
The fourth phase (i.e. Lemma 3) shows that if $p_{n_{k},k+c}\leqslant 2^{c+3}p_{n_{k},k+c+1}$, for $c\in[-k:4]$, then the same is true if we substitute $n_{k}$ with a bigger number (i.e., this property is increasing with respect to the $n$ parameter). In order to apply this result, we have to satisfy some starting conditions. We have numerically checked the appropriate condition for $k=7$ (presented later, in Table 1). Therefore, the second phase of the proof of Theorem 1 justifies the assumptions of Lemma 3 with respect to the parameter $k$, as long as $k\geqslant 7$ and the third phase let us obtain the appropriate assumptions with respect to $c$. One can check that for $k<7$, an analogous assumption is not true.
@@
Let us begin with a reminder. First, $M_{n}$ is ranged in $\mathbb{N}_{0}$ and, moreover, in $I_{n}\subseteq\left[\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4\right]$. We provide few facts about the concentration of the distribution of the random variable $M_{n}$, or more precisely about the probability that $M_{n}$ will be outside the interval $I_{n}$.
@@
Let us mention this with the notions $q_{k}$ and $Q$ from [24], $r_{k}=\frac{1}{q_{k}}$ and $R=\frac{1}{Q}$.
Proof 1
@@
Proof 2
Consider a process $X=(X_{k},k\in\![0:n])$. Let $X$ initially follow the incrementation rule $\mathbb{P}\left(X_{k}=k+1\right)=1$ for $k\in[0:\lceil\log n\rceil+1]$. Afterwards, let this Markov chain imitate the transition rule of Morris Counter, that is
@@
[…]
Theorem 5
@@
Proof 3
Realise that $\mathbb{P}\left(M_{n}\in[1:n+1]\right)=1$. This observation, together with the lemmas 1 and 2 yields
@@
In this part of the investigation, we try to establish the $\varepsilon(n)$ parameter of DP of $M_{n}$. In fact, it remains to examine the property (1) in the interval $I_{n}$, as Theorem 5 entails (2) for $S_{n}=I_{n}$. Therefore, we are interested in finding the upper bound for maximal privacy loss for any $n\in\mathbb{N}$ and $k\in I_{n}$, namely:
@@
The above claim is the result of a simple application of Lemma 6 from A.
@@
Claim 2
If for any given $n$, there exists an ascending and positive sequence $(\alpha_{i})_{i=1}^{n}$ such that $(\forall\;i\in[1:n])\;p_{n,i}=2^{i}\alpha_{i}p_{n,i+1}$, then there also exists an ascending and positive sequence $(\alpha_{i}^{\prime})_{i=1}^{n+1}$ such that
@@
This claim arises from lemmas 7 and 8 from A.
@@
We use Claim 2 to guarantee starting conditions for the next Lemma 3. However, in order to apply Lemma 3, we will also use Claim 1, which assumes that $n\geqslant 2^{7}+1$. Hence, we would like to gather some information about the distribution of $M_{2^{7}+1}$. More precisely, we are interested in the behaviour of $\theta_{i}={p_{129,i}}/{p_{129,i+1}}$ for $i\leqslant 11$, which we present in Table 1.
@@
[…]
@@
Claims 1 and 2, together with Table 1 enable us to apply Lemma 3 for $n=2^{k}+1$ for any $k\geqslant 7$.
Theorem 6
@@
Let us denote $l=\left\lceil\log(n)\right\rceil$ and $c=k-l\in[-4:4]$. Then Lemma 3 bears $p_{n,k-1}\leqslant~2^{c+3}p_{n,k}$, so
@@
Realise that if $n=2^{l-1}+1$ for some $l\in\mathbb{N}$, then a little adjustment is necessary. Indeed, let now $c-1=k-l\in[-4:4]$, and once again, Lemma 3 provides $p_{n-1,k-1}<~2^{c+2}p_{n-1,k}$. However, it still holds that:
@@
[…]
Theorem 6 only provides $\varepsilon(n)\leqslant-\ln\left(1-{16}/{n}\right)$ for $n>128$ (compare with (10)). However, in Figure 5 we may briefly see that the above inequality is true for smaller numbers of requests $n$ as well.
@@
margin:
Proof 6
(of Theorem 1) Suppose that $S_{n}=I_{n}$ in Fact 1. Then from theorems 5 and 6 we can easily see that $\mathbb{P}\left(M_{n}\notin S_{n}\right)<0.00033$ and
@@
[…]
@@ 8.2 General result on Morris’ Counter privacy (Theorem 2)
Remark that $\lceil c\log(\ln(n))\rceil\geqslant 1$ can be always guaranteed, when the $n$ is big enough.
Proof 7
@@
Now, we are going to prove $\delta_{2}^{\ast}:=\mathbb{P}\left(M_{n}\geqslant\lceil\lg(n)\rceil+\rho+1\right)=O\left(n^{-1}\left(\ln(n)\right)^{-c}\right)$. We use a similar notation and technique as in the proof of Lemma 2, but this time we utilise a discrete discrete two-dimensional simplex $(\rho-1)$:
@@
We couple $(M_{n})_{n}$ with the same process $(X_{n})_{n}$ as in Lemma 2. Roughly speaking, $X_{0}=1$ and $X_{n}$ almost always increments by $1$ until $n=\lceil\log(n)\rceil+1$ and it further follows the same incrementation rule as the Morris counter. Then
@@
In addition, we would like to consider fractions $\frac{p_{n+1,k}}{p_{n,k}}$ for $k\in J_{n}(c)$ as in the proof of Theorem 6. Indeed
@@
We are going to use another formula from [24]. For any $n\in\mathbb{N}$ and $k\in[1:n+1]$,
@@
This shows that $\varepsilon(n)=O\left(\frac{(\log(n))^{2}}{n}\right)$$\qed$
@@ 8.3 Main result for MaxGeo Counter (Theorem 3)
where the approximation is the result of the substitution of $l_{\varepsilon}$ without ceiling. $\qed$
@@ 9 Conclusions and Future Work
In Subsection 4.1, we have mentioned the generalisation of the Morris Counter (for bases $a>1$). Analysis of privacy properties of such variants of Morris Counters and various probabilistic counters presented, for example, in [10], [26] may also be promising directions of further research.
In this paper, we focus on the standard definition of differential privacy. However, there is also an issue of preservation of differential privacy for requests given by a group of $k$ users or one individual sending up to $k$ dependent requests over time. This can be described in the language of the so-called $(\varepsilon,\delta)-k$-DP (see the "group privacy scenario" in [18]) . A group of people may tend to behave in the same manner, so they may send $k$ requests in a row. Especially this ”group” may be represented by a single person colluding with the Adversary. It is worth mentioning that this type of generalisation creates an opportunity to modify probabilistic counters so that each incrementation request executes the update request multiple times to reduce the variance of the rescaled estimator. Intuitively, this extension should be especially efficient in preserving the standard differential privacy property when $\varepsilon(n)=\frac{c}{n}+o(n^{-1})$ (as a parameter of standard differential privacy), because both $c$ and $n$ should scale with $k$ linearly. Hence, the next challenging problem is showing that Morris and MaxGeo Counters satisfy $k$-DP property with similar privacy parameters.
The Morris Counter and the MaxGeo Counter are considered the most popular probabilistic counters. However, the results of this paper shed new light on the properties of the probabilistic counter, in general. There is a possibility to provide analogous differential privacy properties for other probabilistic counters. Moreover, this paper enables the provision of differentially private algorithms for other applications, especially those based on Morris or MaxGeo Counter. For example, in Section 4 we mentioned PCSA and HyperLogLog Counters together with their variances, which can be manually adjusted to the applications. The proper choice of the $m$ parameter implies an exchange of memory usage to improve the accuracy of the estimation. We have mentioned that these counters’ differential privacy parameters can be obtained via Observation 1. However, such a direct result may not be satisfying. Hence a more precise calculation is needed. For instance, Observation 1 may be utilised again with some concentration inequalities.
Acknowledgement

### v7

@@ 8.1 Main Theorem for Morris Counter (Theorem 1)
    $\delta$ phase (Theorem 8.8),
@@
    extrapolation of $\mathcal{P}_{k}^{(c)}\leq 2^{c+3}\mathcal{P}_{k}^{(c+1)}$ property to $N>n_{k}$ (Lemma 8.10),
@@
    $\varepsilon$ phase (Theorem 8.12).
During the first phase, we consider a concentration of Morris Counter in the vicinity of its mean value. More precisely, we show that the Morris Counter after $n$ increment requests takes values in relatively small intervals $I_{n}$ with probability $1-\delta$ (note that then $M_{n}$ satisfies condition (2) for $S_{n}=I_{n}$), where $I_{n}$ is defined as in (4) and $\delta$ is some small constant, which arises from the proof. Note that $I_{n}$ can be interpreted as confidence intervals at level $1-\delta$ (see, e.g. Neyman (1937)). This phase is divided into lemmas 1 and 8.6. The first one uniformly bounds the formula for probabilities given by the Theorem below, due to Flajolet:
Theorem 4 (Proposition 1 from Flajolet (1985))
@@
We sum up the bounds on $p_{n,l}$ to obtain a small upper bound for $\delta_{1}:=\mathbb{P}\left(M_{n}\leqslant\left\lceil\log(n)\right\rceil-5\right)$ (Lemma 1). The same bounds cannot be utilised efficiently in the proof of Lemma 8.6. Instead, it couples $M_{n}$ with a process $X_{n}$, which increases during the first $\left\lceil\log(n)\right\rceil+1$ steps, and then follows the same update rule, so $M_{n}\leqslant X_{n}$ almost surely.
Therefore $\delta_{2}:=\mathbb{P}\left(M_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)\leqslant\mathbb{P}\left(X_{n}\geqslant\left\lceil\log(n)\right\rceil+5\right)$, which is much easier to bound from definition. Note that such a coupling cannot be used in the proof of Lemma 1. The first phase is summarised by Theorem 8.8, i.e., establishes $\delta=\delta_{1}+\delta_{2}$.
In the second phase, we show that $(\mathcal{P}_{k}^{(4)})_{k}$ is descending for large enough $k$ and $(\mathcal{P}_{k}^{(5)})_{k}$ is ascending for big enough $k$ (Lemma A.21). A change of monotonicity is the first obstacle in the proof of the main theorem. The main idea beneath the proof of this fact is to calculate the differences between consecutive elements of the considered ”special” sequences by representing them as the sums via application of Flajolet’s Theorem 4 and realising that usually at most first ten terms of the sums are crucial (on the other hand, taking less than eight terms is rarely sufficient). This is the second issue, which makes the proof so complicated. Let us note that Theorem 4 presents an explicit formula for $\mathbb{P}\left(M_{n}=l\right)$, which (as we may experience in A) is not convenient to analyse. However, it is simple enough to find the values numerically (also note that recursive Definition 3 provides those probabilities easily as well, However, this approach is inefficient in terms of memory and time for a large number of requests $n$). Therefore, by precise analysis, we can finally check some sums numerically and obtain the thesis of Lemma A.21 for $k\geqslant 15$. However, numerically, one can extrapolate it to some smaller $k$ as well (remark that this proof does not work for small values of $k$, since not only ten terms of the aforementioned sums are important). The lemma A.21 can be used directly to show that $\mathcal{P}_{k}^{(4)}\leqslant 2^{7}\mathcal{P}_{k}^{(5)}$ for $k\geqslant 7$ (Claim 1). Let us note that, in fact, this result is not true for $k<7$.
@@
The fourth phase (i.e. Lemma 8.10) shows that if $p_{n_{k},k+c}\leqslant 2^{c+3}p_{n_{k},k+c+1}$, for $c\in[-k:4]$, then the same is true if we substitute $n_{k}$ with a bigger number (i.e., this property is increasing with respect to the $n$ parameter). In order to apply this result, we have to satisfy some starting conditions. We have numerically checked the appropriate condition for $k=7$ (presented later, in Table 1). Therefore, the second phase of the proof of Theorem 1 justifies the assumptions of Lemma 8.10 with respect to the parameter $k$, as long as $k\geqslant 7$ and the third phase let us obtain the appropriate assumptions with respect to $c$. One can check that for $k<7$, an analogous assumption is not true.
@@
Let us begin with a reminder. First, $M_{n}\in\mathbb{N}_{0}$ and, moreover,
$I_{n}\subseteq\left[\left\lceil\log(n)\right\rceil-4:\left\lceil\log(n)\right\rceil+4\right].$
We provide few facts about the concentration of the distribution of the random variable $M_{n}$, or more precisely about the probability that $M_{n}$ will be outside the interval $I_{n}$.
@@
Let us mention this with the notions $q_{k}$ and $Q$ from Flajolet (1985), $r_{k}=\frac{1}{q_{k}}$ and $R=\frac{1}{Q}$.
Proof 8.5.
@@
Proof 8.7.
Consider a process $X=(X_{k\in\![0:n]})$. Let $X$ initially follow the incrementation rule $\mathbb{P}\left(X_{k}=k+1\right)=1$ for $k\in[0:\lceil\log n\rceil+1]$. Afterwards, let this Markov chain imitate the transition rule of Morris Counter, that is
@@
Note that when $m<3$ (that is, when $n<7$), then the above sums are empty, but on the other hand $\lceil\log(n)\rceil+5>n+1$, so the inequality is trivially true.
Theorem 8.8.
@@
Proof 8.9.
Realise that $\mathbb{P}\left(M_{n}\in[1:n+1]\right)=1$. This observation, together with the lemmas 1 and 8.6 yields
@@
In this part of the investigation, we try to establish the $\varepsilon(n)$ parameter of DP of $M_{n}$. In fact, it remains to examine the property (1) in the interval $I_{n}$, as Theorem 8.8 entails (2) for $S_{n}=I_{n}$. Therefore, we are interested in finding the upper bound for maximal privacy loss for any $n\in\mathbb{N}$ and $k\in I_{n}$, namely:
@@
The above claim is the result of a simple application of Lemma A.21 from A.
@@
Claim 2.
If for any given $n$, there exists an ascending and positive sequence $(\alpha_{i})_{i=1}^{n}$ such that
$(\forall\;i\in[1:n])\;p_{n,i}=2^{i}\alpha_{i}p_{n,i+1},$
then there also exists an ascending and positive sequence $(\alpha_{i}^{\prime})_{i=1}^{n+1}$ such that
@@
This claim arises from lemmas A.23 and A.25 from A.
@@
We use Claim 2 to guarantee starting conditions for the next Lemma 8.10. However, in order to apply Lemma 8.10, we will also use Claim 1, which assumes that $n\geqslant 2^{7}+1$. Hence, we would like to gather some information about the distribution of $M_{2^{7}+1}$. More precisely, we are interested in the behaviour of $\theta_{i}={p_{129,i}}/{p_{129,i+1}}$ for $i\leqslant 11$, which we present in Table 1.
@@
If we start with c=k+1c=-k+1, then let us prove inductively the appropriate condition for all $N\geqslant n_{k}$. In addition, the thesis is followed by the induction with respect to $c$.
@@
Claims 1 and 2, together with Table 1 enable us to apply Lemma 8.10 for $n=2^{k}+1$ for any $k\geqslant 7$.
Theorem 8.12.
@@
Let us denote $l=\left\lceil\log(n)\right\rceil$ and $c=k-l\in[-4:4]$. Then Lemma 8.10 bears $p_{n,k-1}\leqslant~2^{c+3}p_{n,k}$, so
@@
Realise that if $n=2^{l-1}+1$ for some $l\in\mathbb{N}$, then a little adjustment is necessary. Indeed, let now $c-1=k-l\in[-4:4]$, and once again, Lemma 8.10 provides $p_{n-1,k-1}<~2^{c+2}p_{n-1,k}$. However, it still holds that:
@@
On the other hand, we have inequalities $p_{n+1,k}>\left(1-2^{-l-c}\right)p_{n,k}$ and $p_{n,k}>\left(1-2^{-l-c}\right)p_{n-1,k}$ for any $c\in[-4:4]$, so both fractions exceed 116/n1-{16}/{n}.
Theorem 8.12 only provides $\varepsilon(n)\leqslant-\ln\left(1-{16}/{n}\right)$ for $n>128$ (compare with (10)). However, in Figure 5 we may briefly see that the above inequality is true for smaller numbers of requests $n$ as well.
@@
Proof 8.14.
(of Theorem 1) Suppose that $S_{n}=I_{n}$ in Fact 1. Then from theorems 8.8 and 8.12 we can easily see that $\mathbb{P}\left(M_{n}\notin S_{n}\right)<0.00033$ and
@@
hence, from Fact 1 we obtain the main result.
@@ 8.2 General result on Morris’ Counter privacy (Theorem 2)
Remark that $\lceil c\log(\ln(n))\rceil\geqslant 1$ can always be guaranteed, when $n$ is large enough.
Proof 8.15.
@@
Now, we are going to prove $\delta_{2}^{\ast}:=\mathbb{P}\left(M_{n}\geqslant\lceil\lg(n)\rceil+\rho+1\right)=O\left(n^{-1}\left(\ln(n)\right)^{-c}\right)$. We use a similar notation and technique as in the proof of Lemma 8.6, but this time we utilise a discrete discrete two-dimensional simplex $(\rho-1)$:
@@
We couple $(M_{n})_{n}$ with the same process $(X_{n})_{n}$ as in Lemma 8.6. Roughly speaking, $X_{0}=1$ and $X_{n}$ almost always increments by $1$ until $n=\lceil\log(n)\rceil+1$ and it further follows the same incrementation rule as the Morris counter. Then
@@
In addition, we would like to consider fractions $\frac{p_{n+1,k}}{p_{n,k}}$ for $k\in J_{n}(c)$ as in the proof of Theorem 8.12. Indeed
@@
We are going to use another formula from Flajolet (1985). For any $n\in\mathbb{N}$ and $k\in[1:n+1]$,
@@
This shows that $\varepsilon(n)=O\left(\frac{(\log(n))^{2}}{n}\right)$.
@@ 8.3 Main result for MaxGeo Counter (Theorem 3)
where the approximation is the result of the substitution of $l_{\varepsilon}$ without ceiling.
@@ 9 Conclusions and Future Work
In Subsection 4.1, we have mentioned the generalisation of the Morris Counter (for bases $a>1$). Analysis of privacy properties of such variants of Morris Counters and various probabilistic counters presented, for example, in Csűrös (2010), Fuchs et al. (2012) may also be promising directions of further research.
In this paper, we focus on the standard definition of differential privacy. However, there is also an issue of preservation of differential privacy for requests given by a group of $k$ users or one individual sending up to $k$ dependent requests over time. This can be described in the language of the so-called $(\varepsilon,\delta)-k$-DP (see the ”group privacy scenario” in Dwork and Roth (2014)) . A group of people may tend to behave in the same manner, so they may send $k$ requests in a row. Especially this ”group” may be represented by a single person colluding with the Adversary. It is worth mentioning that this type of generalisation creates an opportunity to modify probabilistic counters so that each incrementation request executes the update request multiple times to reduce the variance of the rescaled estimator. Intuitively, this extension should be especially efficient in preserving the standard differential privacy property when $\varepsilon(n)=\frac{c}{n}+o(n^{-1})$ (as a parameter of standard differential privacy), because both $c$ and $n$ should scale with $k$ linearly. Hence, the next challenging problem is to show that Morris and MaxGeo Counters satisfy the $k$-DP property with similar privacy parameters.
The Morris Counter and the MaxGeo Counter are considered the most popular probabilistic counters. However, the results of this paper shed new light on the properties of the probabilistic counter, in general. There is a possibility to provide analogous differential privacy properties for other probabilistic counters. Moreover, this paper enables the provision of differentially private algorithms for other applications, especially those based on Morris or MaxGeo Counter. For example, in Section 4 we mentioned PCSA and HyperLogLog Counters together with their variances, which can be manually adjusted to the applications. The proper choice of the $m$ parameter implies an exchange of memory usage to improve the accuracy of the estimation. We have mentioned that these counters’ differential privacy parameters can be obtained via Observation 1. However, such a direct result may not be satisfying. Hence a more precise calculation is needed. For example, Observation 1 may be used again with some concentration inequalities.
Acknowledgements.